# On Exponential Instability of an Inverse Problem for the Wave Equation

### v1

@@ 1. Introduction
Let $r\in(0,1)$ be fixed as above, $\Sigma\Subset B_{r}(2e_{1}),\Xi\Subset B_{r}(-2e_{1})$ be open, $T^{\prime}\in(0,T)$, and define $\Omega\coloneqq(0,T^{\prime})\times\Xi$. Let $\mu\in\mathbb{R},\delta>0$ be fixed so that $\mu+\delta>n/2$ and define $K\coloneqq\left\{q\in H^{\mu}(U)\colon\mathrm{supp}\,q\subset\bar{\Sigma}\,,\left\lVert q\right\rVert_{H^{\mu+\delta}(U)}\leq 1\right\}$.
@@
The set $U$ was chosen so that for sources $f\in L^{2}(X_{+})$ with $\mathrm{supp}\,f\subset[0,T]\times\overline{B_{3}(0)}$, by the finite speed of propagation, $S_{q}(f)$ vanishes identically near $(-T,T)\times\partial B_{T+4}(0)$, the ‘far’ boundary of $X$. Thus, the extension of $S_{q}(f)$ by 00 outside of $U$ solves
@@ 2. The Proof
Let $p(t,x,\tau,\xi)=-\tau^{2}+|\xi|^{2}$ be the principal symbol of $\Box$, which will remain fixed throughout. We begin with two purely geometric statements that can be summarized informally as: generalized bicharacteristic arcs of $p$ with a point in $T^{\ast}((0,T)\times B_{r}(-2e_{1}))$ cannot have ‘originated’ from a neighborhood of $T^{\ast}(\{0\}\times B_{r}(2e_{1}))$. The following two statements are immediate when one draws a picture (see Fig. 1), but we give the full details nevertheless, first fixing some terminology. If $H$ is a hyperplane in $\mathbb{R}^{n}$ defined by $H=\{x\colon(x-x^{\circ})\cdot x^{\circ}=0\}$ where $x^{\circ}\in\mathbb{R}^{n}$ is some point, we say that a point $x\in\mathbb{R}^{n}$ lies above (resp. below) $H$ if $(x-x^{\circ})\cdot x^{\circ}>0$ (resp. $<0$).
@@
$0\leq(x_{0}-x^{\circ})\cdot x^{\circ}=-2x^{\circ}_{1}-1+y_{0}\cdot x^{\circ}\leq-2x^{\circ}_{1}-1/4\,,$
@@
$(x-x^{\circ})\cdot x^{\circ}<2x_{1}^{\circ}r-1<r-1\,,$
@@
For every $\nu^{\ast}=(t^{\ast},x^{\ast},\tau^{\ast},\xi^{\ast})\in T^{\ast}((0,T)\times B_{r}(-2e_{1}))\setminus\{0\}$ with $\left\lvert\tau^{\ast}\right\rvert=1$ and $\nu^{\ast}\in p^{-1}(0)$, the forward generalized bicharacteristic arc $\gamma$ (according to [17, Def. 24.3.7] or [25]) on ×$\mathbb{R}\times U$ with initial condition $\gamma(t^{\ast})=\nu^{\ast}$ satisfies $\gamma(t)\in T^{\ast}(\{t\}\times(U\setminus B_{r}(2e_{1})))\setminus\{0\}$ for $t\in[0,t^{\ast}]$. Here, forward means that the projection of $\gamma$ onto the $t$ component is an increasing function.
@@
The following general remark will be used throughout the rest of the proof. Because $\gamma$ travels with unit speed forward in time ($\left\lvert\pi_{\tau}\gamma(t)\right\rvert=\left\lvert\tau^{\ast}\right\rvert=1,t\in[0,t^{\ast}]$) and the ‘far’ boundary ×$\mathbb{R}\times\partial B_{T+4}(0)$ was chosen far away, and $\pi_{x}\gamma(t^{\ast})\in B_{r}(-2e_{1})$, we have $\pi_{x}\gamma(t)\not\in\partial B_{T+4}(0)$ for all $t\in[0,t^{\ast}]$. We make a case distinction, see also Fig. 1 for a pictorial view of the behavior of the ray $\pi_{x}\gamma$.
$B_{r}(-2e_{1})$$B_{r}(2e_{1})$$H$$\mathcal{O}$
@@
Case 2: there is $t^{\circ}\in(0,t^{\ast})$ so that $\pi_{(t,x)}\gamma(t^{\circ})\in(0,T)\times\partial\mathcal{O}$. We remark here that the normal of ×$\mathbb{R}\times\partial\mathcal{O}$ at any $(t,x)\in(0,T)\times\partial\mathcal{O}$ is given by $(0,x)$.
@@
We conclude that $\pi_{(t,x)}\gamma(t)$ lies above the hypersurface $H$ for $t$ near but unequal to $t^{\circ}$. Further (using that $(0,T)\times\mathcal{O}$ lies below $H$ aside from at ×$\mathbb{R}\times\{x^{\circ}\}$), we see that $\pi_{x}\gamma(t)$ does not intersect $\partial U$ for $t\in[0,t^{\ast}]\setminus t^{\circ}$, so that $\pi_{(t,x)}\gamma$ always lies above the hypersurface $H$ for $t\in[0,t^{\ast}]\setminus t^{\circ}$.
@@
Since we will appeal to [20, Thm. 4.2(b)] in order to prove Theorem 1.1, we will have to replace the domain ${Y}$ by some closed manifold. We will also have to show mapping properties of $S_{q}$ into some Banach space rather than $G^{3}$. We thus recall the definition of Gevrey spaces on closed manifolds from [20, § 2.6] and their decomposition into a union of Banach spaces.
@@
Let $d\in\mathbb{N},\sigma\in[1,\infty)$. For a closed smooth manifold $(M,g)$ of dimension $d$ and any $\rho>0$, for $(\varphi_{j})_{j\in\mathbb{N}}\subset L^{2}(M)$ an orthonormal basis of eigenfunctions for $-\Delta_{g}$, we introduce the subspace $A^{\sigma,\rho}(M)$ of $L^{2}(M)$ defined as those $u\in L^{2}(M)$ so that
$\left\lVert u\right\rVert_{A^{\sigma,\rho}(M)}=\left(\sum_{j=0}^{\infty}e^{2\rho j^{\frac{1}{d\rho}}}\left\lvert\langle u,\varphi_{j}\rangle\right\rvert^{2}\right)^{1/2}<\infty\,.$
@@
with the convention that $10^{0}=$ 1.
@@
(2.2) $\left\lVert(-\Delta_{g})^{t}u\right\rVert_{L^{2}(M)}\leq CR^{2t}(2t)^{2t\sigma}$
@@
Recall that we defined ${Y}=(-T,T)\times B_{r}(-2e_{1})$. We set up a functional analytic result that will allow us to replace ${Y}$ by some closed manifold $N\supset{Y}$. It will also later be used to show that $S_{q}-S_{0}$ maps into $A^{3,\rho}$ for some $\rho>0$. We will rely on [21], where for any $\sigma>1$, our notation $G^{\sigma}$ corresponds to $\mathcal{E}^{\{M_{p}\}}$ in that paper, where $M_{p}$ is the sequence $M_{0}=1,M_{p}=p!^{\sigma}\,,p\in\mathbb{N}$.
@@
Consider the torus $N\subset\mathbb{R}^{1+n}$ defined by identifying opposite sides of the cube $[-R,R]^{1+n}\subset\mathbb{R}^{1+n}$, where $R>0$ is chosen sufficiently large so that ${Y}=(-T,T)\times B_{r}(-2e_{1})\Subset N$ (see also [28, Lem. 3.1.8]). With the euclidean metric, $N$ is a closed analytic manifold. We take $\chi\in G^{3}(\mathbb{R}^{1+n})$ to be identically $1$ in an open neighborhood of $\Omega$ and supported in a compact subset of ${Y}\subset N$ (that Gevery cut-offs exist follows from [31, § 1.4] or [16, § 1.4]).
For any $g\in L^{2}(X_{+})$ and $\tilde{q}\in L^{\infty}(U)$, $(S_{\tilde{q}}-S_{0})(g)$ is the unique solution of Eq. 1.1 for $f=-\tilde{q}S_{\tilde{q}}(g)$ and $q=0$, whereas $S_{0}(-\tilde{q}S_{\tilde{q}}(g))$ is also a solution of Eq. 1.1 for the same $f$ and $q$, so that Lemma A.1 guarantees that $S_{0}(-\tilde{q}S_{\tilde{q}}(g))=(S_{\tilde{q}}-S_{0})(g)$. In particular, for any $q\in L^{\infty}(U)$, we have
@@
From Lemma A.1 and Proposition 2.4, we know that $S_{0}\circ\mathbb{1}_{(0,T)\times\Sigma}\colon L^{2}(X_{+})\to H^{1}({Y})$ is continuous and maps $S_{0}\circ\mathbb{1}_{(0,T)\times\Sigma}\colon L^{2}(X_{+})\to G^{3}({Y})$, so that Proposition 2.7 implies that there is $\rho>0$ so that the map
@@
for $b$ from Eq. A.1 and some $C^{\prime}>0$ independent of $q$, which gives Eq. 2.3. ∎
@@
We set up notation to apply [20, Thm. 4.2(b)]. For $N,F_{q}$ from Lemma 2.8, define the operator
@@ A. Existence of Solutions
Translating the condition $u|_{t<0}=0$ in Eq. 1.1 into vanishing Cauchy data at $t=0$, the uniqueness and existence of the solution $u=u_{q}^{f}$ as well as the continuity of $S_{q}$ are guaranteed by [19, Thm. 2.30] (via extension by 00 to $t<0$ of the solution constructed there).

### v2

@@ 1. Introduction
Let $r\in(0,1)$ be fixed, $\Sigma\Subset B_{r}(2e_{1}),\Xi\Subset B_{r}(-2e_{1})$ be open, $T^{\prime}\in(0,T)$, and define $\Omega\coloneqq(0,T^{\prime})\times\Xi$. Let $\mu\in\mathbb{R},\delta>0$ be fixed so that $\mu+\delta>n/2$ and define $K\coloneqq\left\{q\in H^{\mu+\delta}(U)\colon\mathrm{supp}\,q\subset\bar{\Sigma}\,,\left\lVert q\right\rVert_{H^{\mu+\delta}(U)}\leq 1\right\}$.
@@
The set $U$ was chosen so that for sources $f\in L^{2}(X_{+})$ with $\mathrm{supp}\,f\subset[0,T]\times\overline{B_{3}(0)}$, by the finite speed of propagation, $S_{q}(f)$ vanishes identically near $(-T,T)\times\partial B_{T+4}(0)$, the ‘far’ boundary of $X$. Thus, the extension of $S_{q}(f)$ by 0 outside of $U$ solves
@@ 2. The Proof
Let $p(t,x,\tau,\xi)=-\tau^{2}+|\xi|^{2}$ be the principal symbol of $\Box$, which will remain fixed throughout. We begin with two purely geometric statements that can be summarized informally as: generalized bicharacteristic arcs of $p$ with a point in the cotangent bundle $T^{\ast}((0,T)\times B_{r}(-2e_{1}))$ cannot have ‘originated’ from a neighborhood of $T^{\ast}(\{0\}\times B_{r}(2e_{1}))$. The following two statements are immediate when one draws a picture (see Fig. 1), but we give the full details nevertheless, first fixing some terminology. If $H$ is a hyperplane in $\mathbb{R}^{n}$ defined by $H=\{x\colon(x-x^{\circ})\cdot x^{\circ}=0\}$ where $x^{\circ}\in\mathbb{R}^{n}$ is some point, we say that a point $x\in\mathbb{R}^{n}$ lies above (resp. below) $H$ if $(x-x^{\circ})\cdot x^{\circ}>0$ (resp. <0<0).
@@
$0\leq(x_{0}-x^{\circ})\cdot x^{\circ}=-2x^{\circ}_{1}-1+y_{0}\cdot x^{\circ}\leq-2x^{\circ}_{1}-1+r\,$,
@@
$(x-x^{\circ})\cdot x^{\circ}<2x_{1}^{\circ}+r-1<r-1\,$,
@@
For every $\nu^{\ast}=(t^{\ast},x^{\ast},\tau^{\ast},\xi^{\ast})\in T^{\ast}((0,T)\times B_{r}(-2e_{1}))\setminus\{0\}$ with $\left\lvert\tau^{\ast}\right\rvert=1$ and $\nu^{\ast}\in p^{-1}(0)$, the forward generalized bicharacteristic arc $\gamma$ (according to [17, Def. 24.3.7] or [25]) on $\mathbb{R}\times U$ with initial condition $\gamma(t^{\ast})=\nu^{\ast}$ satisfies $\gamma(t)\in T^{\ast}(\{t\}\times(U\setminus B_{r}(2e_{1})))\setminus\{0\}$ for $t\in[0,t^{\ast}]$. Here, forward means that the projection of $\gamma$ onto the $t$ component is an increasing function.
@@
The following general remark will be used throughout the rest of the proof. Because $\gamma$ travels with unit speed forward in time ($\left\lvert\pi_{\tau}\gamma(t)\right\rvert=\left\lvert\tau^{\ast}\right\rvert=1,t\in[0,t^{\ast}]$) and the ‘far’ boundary $\mathbb{R}\times\partial B_{T+4}(0)$ was chosen far away, and $\pi_{x}\gamma(t^{\ast})\in B_{r}(-2e_{1})$, we have $\pi_{x}\gamma(t)\not\in\partial B_{T+4}(0)$ for all $t\in[0,t^{\ast}]$. We make a case distinction, see also Fig. 1 for a pictorial view of the behavior of the ray $\pi_{x}\gamma$.
[…]
@@
Case 2: there is $t^{\circ}\in(0,t^{\ast})$ so that $\pi_{(t,x)}\gamma(t^{\circ})\in(0,T)\times\partial\mathcal{O}$. We remark here that the normal of $\mathbb{R}\times\partial\mathcal{O}$ at any $(t,x)\in(0,T)\times\partial\mathcal{O}$ is given by $(0,x)$.
@@
We conclude that $\pi_{(t,x)}\gamma(t)$ lies above the hypersurface $H$ for $t$ near but unequal to $t^{\circ}$. Further (using that $(0,T)\times\mathcal{O}$ lies below $H$ aside from at $\mathbb{R}\times\{x^{\circ}\}$), we see that $\pi_{x}\gamma(t)$ does not intersect $\partial U$ for $t\in[0,t^{\ast}]\setminus t^{\circ}$, so that $\pi_{(t,x)}\gamma$ always lies above the hypersurface $H$ for $t\in[0,t^{\ast}]\setminus t^{\circ}$.
@@
Since we will appeal to [20, Thm. 4.2(b)] in order to prove Theorem 1.1, we will have to replace the domain Y{Y} by some closed manifold. We will also have to show mapping properties of $S_{q}$ into some Banach space rather than $G^{3}$. We thus recall the definition of Gevrey spaces on closed manifolds from [20, § 2.6] and their decomposition into a union of Banach spaces.
@@
Let $d\in\mathbb{N},\sigma\in[1,\infty)$. For a closed smooth manifold $(M,g)$ of dimension $d$ and any $\rho>0$, for $(\varphi_{j})_{j\in\mathbb{N}}\subset L^{2}(M)$ an orthonormal basis of eigenfunctions for $\Delta_{g}$, we introduce the subspace $A^{\sigma,\rho}(M)$ of $L^{2}(M)$ defined as those $u\in L^{2}(M)$ so that
$\left\lVert u\right\rVert_{A^{\sigma,\rho}(M)}=\left(\sum_{j=0}^{\infty}e^{2\rho j^{\frac{1}{d\sigma}}}\left\lvert\langle u,\varphi_{j}\rangle\right\rvert^{2}\right)^{1/2}<\infty\,.$
@@
with the convention that $0^{0}=1$.
@@
(2.2) $\left\lVert(-\Delta_{g})^{t}u\right\rVert_{L^{2}(M)}\leq CR^{2t}(2t)^{2t\sigma}\$
@@
Recall that we defined ${Y}=(-T,T)\times B_{r}(-2e_{1})$. We set up a functional analytic result that will allow us to replace Y{Y} by some closed manifold $N\supset{Y}$. It will also later be used to show that $S_{q}-S_{0}$ maps into $A^{3,\rho}$ for some $\rho>0$. We will rely on [21], where for any $\sigma>1$, our notation $G^{\sigma}$ corresponds to $\mathcal{E}^{\{M_{p}\}}$ in that paper, where $M_{p}$ is the sequence $M_{0}=1,M_{p}=p!^{\sigma}\,,p\in\mathbb{N}$.
@@
Consider the torus $N\subset\mathbb{R}^{1+n}$ defined by identifying opposite sides of the cube $R,R]^{1+n}\subset\mathbb{R}^{1+n}$, where $R>0$ is chosen sufficiently large so that ${Y}=(-T,T)\times B_{r}(-2e_{1})\Subset N$ (see also [28, Lem. 3.1.8]). With the Euclidean metric, $N$ is a closed analytic manifold. We take $\chi\in G^{3}(\mathbb{R}^{1+n})$ to be identically $1$ in an open neighborhood of $\Omega$ and supported in a compact subset of ${Y}\subset N$ (that Gevrey cut-offs exist follows from [31, § 1.4] or [16, § 1.4]).
For any $g\in L^{2}(X_{+})$ and $\tilde{q}\in L^{\infty}(U)$, $(S_{\tilde{q}}-S_{0})(g)$ is the unique solution of Eq. 1.1 for $f=-\tilde{q}S_{\tilde{q}}(g)$ and $q=0$, whereas $S_{0}(-\tilde{q}S_{\tilde{q}}(g))$ is also a solution of Eq. 1.1 for the same $f$ and $q$, so that Lemma A.1 from the appendix guarantees that $S_{0}(-\tilde{q}S_{\tilde{q}}(g))=(S_{\tilde{q}}-S_{0})(g)$. In particular, for any $q\in L^{\infty}(U)$, we have
@@
From Lemma A.1 (see the appendix) and Proposition 2.4, we know that $S_{0}\circ\mathbb{1}_{(0,T)\times\Sigma}\colon L^{2}(X_{+})\to H^{1}({Y})$ is continuous and maps $S_{0}\circ\mathbb{1}_{(0,T)\times\Sigma}\colon L^{2}(X_{+})\to G^{3}({Y})$, so that Proposition 2.7 implies that there is $\rho>0$ so that the map
@@
for $b$ from the appendix (Equation (A.1)) and some $C^{\prime}>0$ independent of $q$, which gives Eq. 2.3. ∎
@@
We set up notation to apply [20, Thm. 4.2(b)] (or rather the analogous statement thereof when one uses [20, Prop. 3.11] in its proof to allow restricting the support of elements in $K$ as we have done here). For $N,F_{q}$ from Lemma 2.8, define the operator
@@ A. Existence of Solutions
Translating the condition $u|_{t<0}=0$ in Eq. 1.1 into vanishing Cauchy data at $t=0$, the uniqueness and existence of the solution $u=u_{q}^{f}$ as well as the continuity of $S_{q}$ are guaranteed by [19, Thm. 2.30] (via extension by 0 to $t<0$ of the solution constructed there).